Susan and Stan Britton are a married couple who file a joint income tax return, where the tax rates are based on the tax table 3.5. Assume that their taxable income this year was $280,000. Round your answers to two decimal places. What is their federal tax liability? $ What is their marginal tax rate? % What is their average tax rate? %
Answer #1
Tax calculations are based on the rates for 2018.
1) The federal tax liability would be:
=28179+(280000-165000)*24% = $             55,779
2) As the rate is 24% for the slab 165001 to 315000, the marginal tax rate = 24.00%
3) Average rate of tax = Total tax liability/Taxable income = 55779/280000 = 19.92%